OpenStudy (anonymous):

Solve 3x^2-5x-13=0 Type exact answers, using radical as needed. Express complex numbers in term of i. Use a comma to separate answers as needed. f(x)=3x^2-5x-13 What is the x intercept?

OpenStudy (anonymous):

quadratic formula, plug it into that.

OpenStudy (anonymous):

ok I am not sure for the i?

OpenStudy (anonymous):

Radar can u help please

OpenStudy (radar):

I don't think you have an "i" in this problem. Using the quadratic formula,a=3, b=-5 and c=-13 The equation would look like this:\[5\pm \sqrt{5^{2}+156}\over6\]. Note the number in the radical would remain positive. Therefore, no complex number. there is one positive root (3.07) and one negative root -1.41

OpenStudy (anonymous):

So the answers would be 3.07, -1.41

OpenStudy (anonymous):

and the x intercept would be 3.07

OpenStudy (radar):

Yes and they are real you might want to double check these two values of x in your calculator:\[1/6(5-\sqrt{181)}\]\[1/6(5+\sqrt{181})\] for the actual values.

OpenStudy (anonymous):

so that the above answer is what i need to put in there

OpenStudy (anonymous):

So i need to put 6( 5√-181), 6(5√181)

OpenStudy (radar):

The x intercept would be both values, one negative and one positive (there are two intercepts it is a parabola that opens upward and extends below the x axis, where it crosses the x axis are those two points.

OpenStudy (anonymous):

6-√181 i meant

OpenStudy (anonymous):

6(5-√181) i meant

OpenStudy (radar):

That would be the most accurate way to express it, of course if you have a calculator you could express it as an actual decimal number. Let me check it with my calculator.

OpenStudy (anonymous):

So what is the x intercept then please i thought that it was 3.07

OpenStudy (radar):

Those numbers were close enough lol, but the expression with the radicals is totally accurate. The x intercepts are x=-1.409 and x=3.07

OpenStudy (radar):

Note there are two places where this function crosses the x axis (intercepts).

OpenStudy (radar):

if you use the answer with the radicals they are \[5+\sqrt{181}\over 6\] and\[5-\sqrt{181}\over 6\] Note the 5 is part of the answer. In your previous post you removed them for some reason.

OpenStudy (anonymous):

so put them in fraction form

OpenStudy (radar):

What does your teacher want? Are their instructions on the preferred answer form?

OpenStudy (radar):

I personally prefer the answer with the radical and in fraction form just like my very last way I posted it.

OpenStudy (anonymous):

Type exact answers, using radical as needed. Express complex numbers in term of i. Use a comma to separate answers as needed.

OpenStudy (radar):

Remember there are two answers. lol O.K. the fraction with the radical is the way to go.

OpenStudy (anonymous):

Will it be right 6(5-√181), 6(5√181)

OpenStudy (anonymous):

ok I think that I have it thanks I have more to get help on

OpenStudy (radar):

It has to be 1/6 not 6 like this:\[1/6(5+\sqrt{181})\] and the other one just like it except a minus sign between the 5 and radical.

OpenStudy (radar):

Good luck with it collegegirl.

OpenStudy (anonymous):

well I put 5+√181 5-√181 -------- , ------- 6 6

OpenStudy (radar):

Excellent